# ROOFS: RObust biOmarker Feature Selection

### v1

@@ 2.2.6 Automated reporting
To allow users to efficiently evaluate the results of benchmarking, roofs generates a report summarizing the key metrics of FS performance. The report includes: 1) a table that aggregates FS frequencies across methods and shows agreement between different methods in selecting individual features; 2) a stability – AUC plot for a quick visual overview of performance and identification of Pareto-optimal methods; 3) a summary table that combines all metrics from every model, with filtering options that help identify the optimal methods based on user’s criteria of interest.
@@ 2.2.8 Experimental setup
All experiments on PIONeeR data were conducted on a MacBook Pro (2019) with a 2.4 GHz 8-core Intel Core i9 processor and 64 GB 2667 MHz DDR4 RAM.
@@ 3 Results
Figure 2: Instability of LASSO. A: Bootstrap selection frequencies for features selected by LASSO in at least 1 bootstrap sample. B: Variability in selected subset sizes and out-of-bag AUC.
@@ 3.1 LASSO-derived signature shows modest stability
Our motivation in the PIONeeR study was to establish a biomarker signature capable of explaining resistance to immunotherapy in advanced lung cancer. We first applied our benchmarking pipeline, as described in Methods and illustrated in Figure 1, using classical LASSO (Least Absolute Shrinkage and Selection Operator) as a baseline FS method. LASSO remains a standard FS approach in biomedical data analysis due to its simplicity of interpretation and computational efficiency, but it has been shown to be inconsistent (i.e., unable to recover a true model) in the presence of linear dependencies among features, which is common in high-dimensional problems [48]. On the PIONeeR data, LASSO showed moderate stability (S = 0.34, Figure 2A) and high variability in features selected across bootstrap datasets, with 255 out of 374 features selected in at least 1 bootstrap (Figure 2A). LASSO-selected feature sets varied substantially in size, ranging from 13 to 42 predictors, and the corresponding post-selection models, trained using classical ML classifiers, achieved out-of-bag AUC values ($\theta_{\text{out}}$ in the Methods) between 0.49 and 0.78. The apparent LASSO model selected 25 features, of which only 12 were selected with high confidence (defined here as being chosen in at least 50% of bootstrap models), and none were selected across all bootstraps. Low selection frequency of the remaining 13 features raised concerns about their potential clinical relevance and generalizability of the signature.
@@ 3.2 Iterative VIF increases FS stability
p.adjust t-test + logistic .05 0.72 GB 0.39 0.10 ± 0.03 0.93 ± 0.06 1.95 ± 0.09
@@
Results are reported for the optimal FS method-classifier pair, representing the best performance across all tested classifiers. Values are presented as point estimates or mean ± SD (standard deviation); run times correspond to the average duration of the FS step (excluding predictive modeling) across bootstraps. Better than LASSO. GB: Gradient boosting classifier; St: stacking classifier; Ex: extra trees classifier; Bag: bagging classifier; LR: logistic regression; SVC: support vector classifier; LDA: linear discriminant analysis; MLP: multi-layer perceptron; LASSO: least absolute shrinkage and selection operator; HSIC: Hilbert-Schmidt independence criterion; RENT: repeated elastic net technique; CIFE: conditional infomax feature extraction; CMIM: conditional mutual information maximization; DISR: double input symmetrical relevance; JMI: joint mutual information; mRMR: minimum Redundancy Maximum Relevance; LR-RFE: logistic regression-based recursive feature elimination.
@@ 3.3 Comprehensive benchmarking shows variability in FS stability and performance
On the reduced post-VIF dataset, filters were generally the most stable and predictive family of FS methods, although variability was observed within the group. Specifically, statistical filters achieved high optimism-corrected AUC values (ranging from 0.7 to 0.72), and several of them demonstrated good stability, for example, Fisher score (S = 0.37), t-score (S = 0.36), and p.adjust (S = 0.39) (see Table 2). In contrast, CIFE and DISR, filters based on information theory scores, achieved top stability values (S = 0.73 and 0.54, respectively) but lower predictive performance (optimism-corrected AUC = 0.67 and 0.61, respectively). All information theory filters (CIFE, CMIM, DISR, JMI) often omitted known biomarkers (e.g. lactate dehydrogenase and C-reactive protein), showing that selection can be consistent but wrong. CMIM and JMI were both unstable and not predictive, although still outperforming random FS. Embedded methods exhibited the lowest stability among other groups at comparable predictive performance. Surprisingly, embedded resampling-based ensembles designed to improve the consistency of FS (stability selection, Stabl LASSO, and RENT) did not outperform standard LASSO neither in stability nor in AUC. Their stringent selection may account for their lower predictive performance, but it does not explain the lack of improvement in stability (S = 0.3, 0.25, and 0.29, respectively, vs LASSO S = 0.34). The instability of these methods could be attributed to the variability behind the resampling procedures on which they rely, although this requires further investigation. The performance of the wrapper methods was intermediate, with AUC (0.69-0.70) and stability (0.32-0.38) falling between filter methods (the best) and embedded methods (the worst).
@@
Among all methods, the p.adjust, a statistical filter that combined t-test and logistic regression adjusted p-values at a 0.05 threshold, achieved the best balance between stability (S = 0.39) and performance (AUC = 0.72) (Figure 4A). The gradient boosting (GB) model trained on the 22-feature signature selected using the p.adjust achieved predictive performance equivalent to that of the GB model trained on the full feature set (214 features) without FS (AUC = 0.71). Note that the strong performance of the no-FS GB model can be attributed to the restriction of the depth of tree-based classifiers in the benchmark to 2, which added an additional internal layer of FS.
@@
Overall, most benchmarked methods achieved optimism-corrected AUC values comparable to standard LASSO, although several approaches, predominantly statistical filters, showed modest improvements. However, more than half of the methods demonstrated higher stability than LASSO, supporting the idea that it is possible to select features more consistently while maintaining similar predictive performance [36]. Across all methods, the association between stability and AUC was weak (Spearman’s $r=0.32$, $p=0.044$). The least stable approaches tended to perform poorly, but high stability did not necessarily translate into top predictive performances.
@@
Figure 4: Performance of representative FS methods from different algorithmic families (top: FS methods with fixed, user-defined subset size; bottom: FS methods with subset size based on algorithm thresholds). Blue: embedded methods; violet-blue: resampling-based ensemble methods; green: filters; rose: wrappers. A, C: Trade-off between discovery and error in the experiment on the semi-synthetic PIONeeR dataset with a simulated outcome. The true positive rate denotes the proportion of true predictors correctly identified by each FS method, while the false positive rate shows the proportion of non-informative features incorrectly selected. Error bars represent one-half of the standard deviation (SD). B, D: Stability and predictive performance (AUC) in the experiment on the real PIONeeR dataset.
@@
Even though the outcome was generated with a simple linear model, no method achieved both high TPR and low FPR (Table 2, fig. 4B). By construction, fixed-size methods exhibited an inverse relationship between TPR and FPR, as true discoveries constrained the number of false positives (fig. 4). In contrast, threshold-based methods displayed the opposite behavior: methods with higher TPR also had higher FPR. More relaxed methods (e.g., LASSO) showed higher TP and FP rates, whereas more stringent methods (e.g., Stable LASSO) produced smaller TP and FP rates. The same behavior was observed when FPs were represented as the false discovery rate (FDR), i.e., the proportion of FPs relative to the signature size (supplementary fig. S1).
@@
As expected, embedded ensemble methods that control for false discoveries (stability selection, Stabl LASSO, RENT) achieved the lowest FPR but at the cost of lower TPR. These methods are characterized by consistent strict selection of features with the strongest signal. However, they omit a high proportion of true predictors with weaker signals, demonstrating the TPR of only 0.46 ± 0.07, 0.31 ± 0.07 and 0.47 ± 0.07, respectively. Statistical filters exhibited the highest TPR values, while their FPR remained moderate, with one exception. p.adjust filter achieved the highest TPR (0.93 ± 0.06), rarely omitting true predictors, but at the cost of high FPR and FDR, despite BH adjustment being explicitly designed to control false discoveries. FPR of p.adjust was significantly higher than that of random FS (0.1 ± 0.03 vs 0.07 ± 0.03; Mann–Whitney U test, $p<0.0001$), while its FDR was high but significantly lower than that of random FS (0.53 ± 0.06 vs 0.89 ± 0.04; Mann–Whitney U test, $p<0.0001$). The high FPR of p.adjust may be explained by residual correlations among predictors in the post-VIF dataset, which can inflate false discoveries in multiple hypothesis testing settings, as demonstrated by Kanduri et al. [29]. Finally, information theory filters performed poorly, exhibiting both high FPR and low TPR, consistent with their omission of known biomarkers in the real outcome analysis.
@@
Overall, Table 2 illustrates that a low FPR did not necessarily correspond to better predictive performance, whereas higher TPR values were generally associated with improved AUC. This highlights that generalisation requires identifying a sufficient set of true predictors rather than minimising false positives alone.
@@
Table 2 additionally compares computational efficiency across FS methods. Most filter and embedded methods required $\leq 10$ seconds per bootstrap iteration (and $\leq 1$ second for some). In contrast, wrapper methods and resampling-based approaches were substantially more demanding, with computation time approaching 1 minute per bootstrap iteration.
3.5 p.adjust filter achieves a more stable and predictive signature
@@
Based on the results presented above, the p.adjust filter based on the union of BH-adjusted p-values from logistic regression (controlled for PD-L1) and t-test with 0.05 threshold was identified as the top-performing method for predicting the resistance to immunotherapy in the PIONeeR data. It demonstrated the highest optimism-corrected AUC (0.72) when fitted with the gradient boosting classifier. It also ranked among the most stable methods (S = 0.39). When benchmarked on semi-synthetic dataset, it achieved the highest TPR of 0.93 ± 0.06 at the expense of a quite high FPR of 0.1 ± 0.03. This was acceptable in our setting, as the cost of missing a true biomarker was considered to be higher than that of including redundant biomarkers. Importantly, unlike standard LASSO and Shapicant, a representative of wrapper methods, all p.adjust-derived features were selected in $50\%$ of bootstrap samples, increasing the reliability of the resulting signature (Figure 5A).
@@
Additionally, we evaluated the stability of the p.adjust models at the individual patient level as measured by the instability index [38] – the proportion of bootstrap models that classify a given patient differently compared to the full model (Figure 5B). Ideally, the instability index should be elevated only for patients whose predicted probabilities lie close to the decision threshold of the full model. All models lacked this behavior. However, overall p.adjust improved patient-level stability: only 23.4% of patients had an instability index exceeding 0.1, compared to 29.9% for Shapicant and 32.0% for LASSO. At the more strict threshold of 0.3, this proportion decreased to 9.4% of patients for p.adjust, versus 12.4% for Shapicant and 9.9% for LASSO.
@@
Although the signature selected by the p.adjust on the full dataset comprised 22 features, bootstrap-derived subsets included up to 72 features (Figure 5C). We investigated this by repeating the experiment using subsampling without replacement at 80% of the original sample size. Under subsampling, the inflation of selected subset size disappeared, while stability further increased and predictive performance was preserved. This suggests that the larger subset sizes observed with bootstrapping may result from over-representation of specific patient subclusters, increasing sample-specific signals that are weak in the full dataset. Therefore the sparsity instability of p.adjust appears to result from the bootstrapping procedure rather than the method itself.
@@
Figure 5: Comparison of performance between Shapicant (representative wrapper method), LASSO (baseline embedded method), and p.adjust (the FS approach selected using roofs). A: Selection frequency of signature features. B: Instability index, defined as the proportion of bootstrap models yielding a different classification for a given patient compared to the full model [38]. C: Out-of-bag AUC ($\theta_{\text{OOB}}$; see Methods) as a function of selected subset size across bootstrap models.
@@
In conclusion, conducting the FS benchmark using roofs and comparing the results obtained by multiple FS methods allowed us to improve performance and reliability from the baseline FS approach by identifying p.adjust as a method with optimal stability and AUC on the data with the real outcome and the best balance between TPR and FPR on the data with the simulated outcome.

### v2

@@ 2.2.6 Automated reporting
To allow users to efficiently evaluate the results of benchmarking, ROOFS generates a report summarizing the key metrics of FS performance. This report includes: 1) a table that aggregates FS frequencies across methods and shows agreement between different methods in selecting individual features; 2) a stability – AUC plot for a quick visual overview of performance and identification of Pareto-optimal methods; 3) a summary table that combines all metrics from every model, with filtering options that help identify the optimal methods based on the user’s criteria of interest. For an example, see the PIONeeR FS report generated with ROOFS: https://compo.inria.fr/pioneer-website/technical_reports/ML_supervised/feature_selection/main_study/
@@ 2.2.8 Experimental setup
All experiments on PIONeeR data were conducted on a MacBook Pro (2019) with a 2.4 GHz 8-core Intel Core i9 processor and 64 GB 2667 MHz DDR4 RAM. The analysis was performed with Python 3.10.2 and R 4.2.2.
@@ 3 Results
Figure 2: Instability of LASSO on the full PIONeeR dataset ($p=374$). A: Bootstrap selection frequencies for features selected by LASSO in at least 1 bootstrap sample. B: Variability in selected subset sizes and out-of-bag AUC.
@@ 3.1 LASSO-derived signature shows modest stability
Motivated by the goal of establishing a biomarker signature capable of explaining resistance to immunotherapy in advanced lung cancer patients from a PIONeeR study, we first applied our benchmarking pipeline, as described in Methods and illustrated in Figure 1, using classical LASSO (Least Absolute Shrinkage and Selection Operator) as a baseline FS method. LASSO remains a standard FS approach in biomedical data analysis due to its simplicity of interpretation and computational efficiency, but it has been shown to be inconsistent (i.e., unable to recover a true model) in the presence of linear dependencies among features, which is common in high-dimensional problems [53]. On the PIONeeR data, LASSO showed moderate stability (S = 0.34, Figure 2A) and considerable variability in features selected across bootstrap datasets, with 255 out of 374 features selected in at least 1 bootstrap (Figure 2A). LASSO-selected feature sets varied substantially in size, ranging from 13 to 42 predictors, and the corresponding post-selection models, trained using classical ML classifiers, achieved out-of-bag AUC values ($\theta_{\text{out}}$ in the Methods) between 0.49 and 0.78. The apparent LASSO model selected 25 features, of which only 12 were selected with high confidence (defined here as being chosen in at least 50% of bootstrap models), and none were selected across all bootstraps. Low selection frequency of the remaining 13 features raised concerns about their potential clinical relevance and generalizability of the signature.
@@ 3.2 Iterative VIF increases FS stability
p.adjust t-test + logistic .05 0.72 GB 0.39 0.10 ± 0.03 0.92 ± 0.06 1.95 ± 0.09
@@
AUC is reported for the optimal FS method-classifier pair across all tested combinations. Values are presented as point estimates or mean ± SD (standard deviation); run times correspond to the average duration of the FS step (excluding predictive modeling) across bootstraps. Corresponds to better than LASSO. GB: Gradient boosting classifier; St: stacking classifier; Ex: extra trees classifier; Bag: bagging classifier; LR: logistic regression; SVC: support vector classifier; LDA: linear discriminant analysis; MLP: multi-layer perceptron; LASSO: least absolute shrinkage and selection operator; HSIC: Hilbert-Schmidt independence criterion; RENT: repeated elastic net technique; CIFE: conditional infomax feature extraction; CMIM: conditional mutual information maximization; DISR: double input symmetrical relevance; JMI: joint mutual information; mRMR: minimum Redundancy Maximum Relevance; LR-RFE: logistic regression-based recursive feature elimination.
@@ 3.3 Comprehensive benchmarking shows variability in FS stability and performance
On the reduced post-VIF dataset, filters were generally the most stable and predictive family of FS methods, although variability was observed within the group. Specifically, statistical filters achieved high optimism-corrected AUC values (ranging from 0.7 to 0.72), and several of them demonstrated good stability, for example, Fisher score (S = 0.37), t-score (S = 0.36), and p.adjust (S = 0.39) (see Table 2). In contrast, CIFE and DISR, filters based on information theory scores, achieved top stability values (S = 0.73 and 0.54, respectively) but lower predictive performance (optimism-corrected AUC = 0.67 and 0.61, respectively). All information theory filters (CIFE, CMIM, DISR, JMI) often omitted known biomarkers (e.g. lactate dehydrogenase and C-reactive protein), showing that selection can be consistent but wrong. CMIM and JMI were both unstable and not predictive, although still outperforming random FS. Embedded methods exhibited the lowest stability among other groups at comparable predictive performance. Surprisingly, embedded resampling-based ensembles designed to improve the consistency of FS (Bolasso, stability selection, Stabl LASSO, RENT, and HSIC LASSO) did not outperform standard LASSO neither in stability nor in AUC. Their stringent selection (with the exception of HSIC LASSO) may account for their lower predictive performance, but it does not explain the lack of improvement in stability (S = 0.34, 0.3, 0.25, 0.29, and 0.25, respectively, vs LASSO S = 0.34). The instability of these methods could be attributed to the variability behind the resampling procedures on which they rely, although this requires further investigation. The performance of the wrapper methods was intermediate, with AUC (0.69-0.70) and stability (0.32-0.38) falling between filter methods (the best) and embedded methods (the worst).
@@
Among all methods, the p.adjust, a statistical filter that combined t-test and logistic regression adjusted p-values at a 0.05 threshold, achieved the best balance between stability (S = 0.39) and performance (AUC = 0.72) (Figure 4A). The gradient boosting (GB) model trained on the 22-feature signature selected using the p.adjust achieved predictive performance equivalent to that of the GB model trained on the full feature set (214 features) without FS (AUC = 0.71). Note that the strong performance of the FS-free GB model can be attributed to the restriction of the depth of tree-based classifiers in the benchmark to 2, which added an additional internal layer of FS.
@@
Overall, most benchmarked methods achieved optimism-corrected AUC values comparable to standard LASSO, with statistical filters and wrappers showing modest improvements and information-theory filters performing worse. However, more than half of the methods demonstrated higher stability than LASSO, supporting the idea that it is possible to select features more consistently while maintaining similar predictive performance [41]. Across all methods, the association between stability and AUC was weak (Spearman’s $r=0.32$, $p=0.044$). The least stable approaches tended to perform poorly, but high stability did not necessarily translate into top predictive performances.
@@
Figure 4: Performance of representative FS methods from different algorithmic families (top: FS methods with fixed, user-defined subset size; bottom: FS methods with varying subset size dependent on algorithm thresholds). Blue: embedded methods; violet-blue: resampling-based ensemble methods; green: filters; rose: wrappers. A, C: Trade-off between discovery and error in the experiment on the semi-synthetic PIONeeR dataset with a simulated outcome. The true positive rate (TPR) denotes the proportion of true predictors correctly identified by each FS method, while the false positive rate (FPR) shows the proportion of non-informative features incorrectly selected. Error bars represent one-half of the standard deviation (SD). B, D: Stability and predictive performance (AUC) in the experiment on the real post-VIF PIONeeR dataset.
@@
Even though the outcome was generated with a simple linear model, no method achieved both high TPR and low FPR (Table 2, fig. 4A, C). By construction, fixed-size methods exhibited an inverse relationship between TPR and FPR, as true discoveries constrained the number of false positives (fig. 4A). In contrast, threshold-based methods displayed the opposite behavior: methods with higher TPR also had higher FPR (fig. 4C). More relaxed methods (e.g., LASSO) showed higher TP and FP rates, whereas more stringent methods (e.g., Stable LASSO) produced lower TP and FP rates. The same behavior was observed when FPs were represented as the false discovery rate (FDR), i.e., the proportion of FPs relative to the signature size (supplementary fig. S1).
@@
As expected, embedded ensemble methods that control for false discoveries (Bolasso, stability selection, Stabl LASSO, RENT) achieved the lowest FPR but at the cost of lower TPR. These methods are characterized by consistent strict selection of features with the strongest signal. However, they omit a high proportion of true predictors with weaker signals, demonstrating the TPR of only 0.49 ± 0.07, 0.46 ± 0.07, 0.31 ± 0.07 and 0.47 ± 0.07, respectively. Statistical filters exhibited the highest TPR values, while their FPR remained moderate, with one exception: p.adjust filter achieved the highest TPR (0.93 ± 0.06), rarely omitting true predictors, but at the cost of high FPR and FDR, despite Benjamini-Hochberg adjustment being explicitly designed to control false discoveries. FPR of p.adjust was significantly higher than that of random FS (0.1 ± 0.03 vs 0.07 ± 0.03; Mann–Whitney U test, $p<0.001$), while its FDR was high but significantly lower than that of random FS (0.53 ± 0.06 vs 0.89 ± 0.04; Mann–Whitney U test, $p<0.001$). The high FPR of p.adjust may be explained by residual correlations among predictors in the post-VIF dataset, which can inflate false discoveries in multiple hypothesis testing settings, as demonstrated by Kanduri et al. [32]. This may be further investigated by applying stricter VIF thresholds, which was not addressed in the present study. Finally, information theory filters performed poorly, exhibiting both high FPR and low TPR, consistent with their omission of known biomarkers in the real outcome analysis.
@@
Overall, a low FPR did not necessarily translate into better predictive performance, while higher TPR values were strongly associated with a higher AUC (Spearman’s $r=0.8$, $p<0.001$). This aligns with the intuitive expectation that strong generalization relies on identifying a sufficiently rich set of true predictors, rather than focusing solely on minimizing false positives.
@@
Table 2 additionally compares computational efficiency across FS methods. Most filter and embedded methods required $\leq 10$ seconds per bootstrap iteration (and $\leq 1$ second for some). Wrapper methods and resampling-based approaches were substantially more demanding, with computation time approaching 1 minute per bootstrap iteration.
3.5 p.adjust filter achieved a more stable and predictive signature
@@
The design of the PIONeeR study prioritised predictive performance, stability, and discovery capability (i.e., high TPR) in the evaluation of FS methods. A higher FPR was considered acceptable, as the cost of missing a true biomarker was greater than that of including redundant biomarkers. These considerations led to the identification of the p.adjust filter using the union of BH-adjusted p-values from logistic regression (adjusted for PD-L1) and the t-test at a 0.05 threshold as the top-performing method for predicting resistance to immunotherapy on PIONeeR data. The p.adjust demonstrated the highest optimism-corrected AUC (0.72) when fitted with the gradient boosting classifier. It also ranked among the most stable methods (S = 0.39). When benchmarked on semi-synthetic dataset, it achieved the highest TPR of 0.93 ± 0.06 at the expense of FPR of 0.1 ± 0.03.
@@
Importantly, unlike standard LASSO, a baseline embedded FS method, and Shapicant, a top-performing representative of wrapper methods, all p.adjust-derived features were selected in $50\%$ of bootstrap samples, increasing the reliability of the resulting signature (Figure 5A). Additionally, we evaluated the stability of the p.adjust models at the individual patient level as measured by the instability index [44] – the proportion of bootstrap models that classify a given patient differently compared to the full model (Figure 5B). Ideally, the instability index should be elevated only for patients whose predicted probabilities lie close to the decision threshold of the full model. All models lacked this behavior. However, overall p.adjust improved patient-level stability: only 23.4% of patients had an instability index exceeding 0.1, compared to 29.9% for Shapicant and 32.0% for LASSO. At the more strict threshold of 0.3, this proportion decreased to 9.4% of patients for p.adjust, versus 12.4% for Shapicant and 9.9% for LASSO.
@@
Although the signature selected by the p.adjust on the full dataset comprised 22 features, bootstrap-derived subsets included up to 72 features (Figure 5C). We investigated this by repeating the experiment using subsampling without replacement at 80% of the original sample size. Under subsampling, the inflation of the selected subset size disappeared, while stability increased and predictive performance was preserved. A similar reduction in selected subset size was observed for the majority of benchmarked methods under resampling without replacement. This suggests that the larger subset sizes observed with bootstrapping may result from over-representation of specific patient subclusters, increasing sample-specific signals that are weak in the full dataset. Therefore the sparsity instability of p.adjust appears to result from the bootstrapping procedure rather than the method itself.
@@
Figure 5: Comparison of performance between Shapicant (representative wrapper method), LASSO (baseline embedded method), and p.adjust (the FS approach selected using ROOFS). A: Selection frequency of signature features. B: Instability index, defined as the proportion of bootstrap models leading to a different classification decision for a given patient compared to the full model [44]. C: Out-of-bag AUC ($\theta_{\text{OOB}}$; see Methods) as a function of selected subset size across bootstrap models.
@@
In conclusion, conducting the FS benchmark using ROOFS and comparing the results obtained by multiple FS methods allowed us to improve signature performance and robustness compared to the baseline FS approach (LASSO) by identifying p.adjust as a method with optimal stability and AUC on the data with the real outcome and the best balance between TPR and FPR on the data with the simulated outcome.